# Emotions Where Art Thou: Understanding and Characterizing the Emotional Latent Space of Large Language Models

### v1

@@ 1 Introduction
The literature on affect in NLP has focused on sentiment analysis, a task where models classify inputs into discrete emotional or affective categories Prabowo & Thelwall (2009); Medhat et al. (2014); Wadawadagi & Pagi (2020); Wankhade et al. (2022). While this demonstrates that LLMs can identify emotions, it offers little insight into how emotional meaning is represented internally. Classification accuracy is not equivalent to interpretability.
Other works have taken a behavioral view, exploring the emotional “intelligence” of LLMs. These include prompting models with hypothetical emotional scenarios and evaluating their responses Wang et al. (2023), or probing how well they align with human judgments in affective tone Huang et al. (2024); Zhao et al. (2023). Though these studies suggest some degree of affective sensitivity, they focus on testing outputs rather than investigating internal mechanisms of the LLM.
Recent work has also examined emotion manipulation and decoding. For instance, models have been used to map text to dimensional emotion ratings like valence-arousal-dominance (VAD) Shah et al. (2023); Broekens et al. (2023), or to generate emotionally inflected language on demand Reichman et al. (2025). LLMs have also been shown to be more likely to comply with emotionally framed requests Vinay et al. (2024). These studies also treat emotion primarily as a label or generation condition—not a latent internal representation.
@@ 2 Related Works
Models of Emotions. Psychological models of emotion are commonly categorized as either discrete or continuous. Discrete theories posit that emotions are fundamentally distinct categories—such as the six “basic” emotions proposed by Ekman et al. (1999): anger, surprise, disgust, enjoyment, fear, and sadness. Other taxonomies expand this set, including more nuanced affective states Plutchik (1991).
In contrast, continuous models view emotions as points in a low-dimensional latent space. A widely used formulation is the valence-arousal-dominance (VAD) model Mehrabian (1996), where valence encodes hedonic tone, arousal measures intensity, and dominance reflects control or agency. Variants of this framework reduce or alter the axes (e.g., Russell’s 2D circumplex Russell (1980)).
@@
Neuroscientific models of emotion offer a parallel debate. Localist theories posit that discrete emotions correspond to specific, anatomically distinct brain regions, while constructionist theories argue that emotions emerge from distributed, domain-general processes Lindquist et al. (2012); Vytal & Hamann (2010); Celeghin et al. (2017). Our results, particularly from ML-AURA (Section 5), support a constructionist-style interpretation in LLMs: emotional content is not localized to a small subset of units but is instead widely distributed across neurons and layers, with high separability emerging from overlapping, multi-purpose components.
Emotions in Latent Space. Recent work has investigated how LLMs interact with emotional text, often focusing on behavior or output-level mappings. For example, ChatGPT has shown the ability to map emotions to Valence-Arousal-Dominance (VAD) values Broekens et al. (2023); Yongsatianchot et al. (2023), suggesting that emotion-relevant dimensions are accessible to the model. However, such studies do not analyze the internal structure or geometry of these latent representations.
Some prior work explicitly trains models to embed emotions into structured spaces, using classification objectives or external supervision. For instance, Dathathri et al. (2019) and Buechel et al. (2020) train models to map between emotion spaces. Similarly, Wang & Zong (2021) learns an emotion space from labeled data, shows clustering by valence, and demonstrates transferability across datasets. However, in all of these works, the emotion space is imposed or supervised, not emergent.
A growing line of work probes how pretrained models encode emotion. Hollinsworth et al. (2024) show that valence is linearly embedded in contextual states, while Zhang et al. (2023) find arousal and dominance less separable, though their analysis depends on encoder-only models and fixed affective lexicons. In contrast, we study decoder-only LLMs, seeking to recover emergent emotional structure directly from hidden-state geometry rather than imposing a psychological model.
Other studies have shown that LLMs exhibit strong zero-shot emotion classification performance across languages Bianchi et al. (2022), though subsequent work notes that language-specific tuning is sometimes necessary for culturally grounded affect De Bruyne et al. (2022). These findings suggest that emotion representations are at least partially transferable across linguistic domains—a hypothesis we test more directly through geometric alignment and projection-based analysis in Section 4.
@@ 3 Methods
To understand how emotions are represented in LLMs, a variety of tools were used. This section outlines those methods and their theoretical grounding. Empirical findings from these analyses are presented in Sections 4 and 5.
ML-AURA. ML-AURA quantifies how selectively a neuron responds to a specific concept by framing each neuron as a threshold-based detector Suau et al. (2024). For a labeled dataset $D$, each neuron’s output is summarized per example using the maximum activation across tokens. These scalar responses are then ranked and evaluated using the area under the precision-recall curve, comparing neuron output against the presence or absence of the target concept. Neurons with high AUC-PR are designated as “experts” for that concept.
In our adaptation, the concepts are emotion categories. We apply ML-AURA in a one-vs-all setup for each emotion, scoring each neuron by how well it distinguishes a target emotion from all others.
@@
Centered-SVD. We build on prior work showing that LLM hidden states lie on low-dimensional manifolds where semantic and syntactic properties are linearly recoverable Aghajanyan et al. (2020); Hu et al. (2022); Lizzo & Heck (2025). To isolate emotion-relevant subspaces, we apply singular value decomposition (SVD) to hidden state activations.
We use the dataset from Reichman et al. (2025), which pairs neutral sentences with synthetic rewrites expressing target emotions. Sentence-level hidden states are obtained by mean-pooling token activations. The resulting activation matrix is centered and decomposed with SVD, yielding principal directions in representation space. Since emotional content is the primary structured variation across inputs, we hypothesize that the leading components capture dominant emotional axes. We test this through alignment, probing, and causal manipulation.
@@
Space Alignment. Prior work has shown that latent spaces arising from related tasks often exhibit similar internal geometry, with relationships between them approximately rigid or linear up to rescaling and rotation Moschella et al. (2023). While some approaches lift these spaces into anchor-relative representations to handle isometric variance, recent work demonstrates that direct alignment via linear or rigid transformations is often sufficient and easier to apply in practice Lähner & Moeller (2024). Following this approach, we use linear regression to align the emotional subspace derived from synthetic data with that derived from human-authored emotion classification datasets. This alignment allows us to test whether the structure found in the synthetic manifold reflects transferable emotional encodings or artifacts specific to the synthetic generation process.
4 Emotion Universality
@@ 4.1 Datasets
We evaluate the universality of emotional representations in LLMs using eight diverse datasets, each offering explicit categorical emotion labels; datasets restricted to polarity or star ratings were excluded as too coarse. The collection spans languages, modalities, and styles: Go-Emotions contains English Reddit comments Demszky et al. (2020), CARER covers English tweets Saravia et al. (2018), SemEval-2007 Task 14 focuses on English news headlines Strapparava & Mihalcea (2007), EmoEvent includes English and Spanish tweets Plaza-del-Arco et al. (2020), Emotions in Drama consists of German plays from the eighteenth and nineteenth centuries Dennerlein et al. (2023), Bhaav offers Hindi short stories Kumar et al. (2019), MultiEmotions-It provides Italian YouTube and Facebook comments Sprugnoli et al. (2020), and EmoTextToKids features French journalistic and encyclopedic texts written for children Étienne et al. (2024). Appendix A analyzes the structural and stylistic contents of each of these datasets.

### v2

@@ 1 Introduction
The literature on affect in NLP has focused on sentiment analysis, a task where models classify inputs into discrete emotional or affective categories (Prabowo and Thelwall, 2009; Medhat et al., 2014; Wadawadagi and Pagi, 2020; Wankhade et al., 2022). While this demonstrates that LLMs can identify emotions, it offers little insight into how emotional meaning is represented internally. Classification accuracy is not equivalent to interpretability.
Other works have taken a behavioral view, exploring the emotional “intelligence” of LLMs. These include prompting models with hypothetical emotional scenarios and evaluating their responses (Wang et al., 2023), or probing how well they align with human judgments in affective tone (Huang et al., 2024; Zhao et al., 2023). Though these studies suggest some degree of affective sensitivity, they focus on testing outputs rather than investigating internal mechanisms of the LLM.
Recent work has also examined emotion manipulation and decoding. For instance, models have been used to map text to dimensional emotion ratings like valence-arousal-dominance (VAD) (Shah et al., 2023; Broekens et al., 2023), or to generate emotionally inflected language on demand (Reichman et al., 2025). LLMs have also been shown to be more likely to comply with emotionally framed requests (Vinay et al., 2024). These studies also treat emotion primarily as a label or generation condition—not a latent internal representation.
@@ 2 Related Works
Models of Emotions. Psychological models of emotion are commonly categorized as either discrete or continuous. Discrete theories posit that emotions are fundamentally distinct categories—such as the six “basic” emotions proposed by (Ekman et al., 1999): anger, surprise, disgust, enjoyment, fear, and sadness. Other taxonomies expand this set, including more nuanced affective states (Plutchik, 1991).
In contrast, continuous models view emotions as points in a low-dimensional latent space. A widely used formulation is the valence-arousal-dominance (VAD) model (Mehrabian, 1996), where valence encodes hedonic tone, arousal measures intensity, and dominance reflects control or agency. Variants of this framework reduce or alter the axes (e.g., Russell’s 2D circumplex (Russell, 1980)).
@@
Neuroscientific models of emotion offer a parallel debate. Localist theories posit that discrete emotions correspond to specific, anatomically distinct brain regions, while constructionist theories argue that emotions emerge from distributed, domain-general processes (Lindquist et al., 2012; Vytal and Hamann, 2010; Celeghin et al., 2017). Our results, particularly from ML-AURA (Section 4.3), support a constructionist-style interpretation in LLMs: emotional content is not localized to a small subset of units but is instead widely distributed across neurons and layers, with high separability emerging from overlapping, multi-purpose components.
Emotions in Latent Space. Recent work has investigated how LLMs interact with emotional text, often focusing on behavior or output-level mappings. For example, ChatGPT has shown the ability to map emotions to Valence-Arousal-Dominance (VAD) values (Broekens et al., 2023; Yongsatianchot et al., 2023), suggesting that emotion-relevant dimensions are accessible to the model. However, such studies do not analyze the internal structure or geometry of these latent representations.
Some prior work explicitly trains models to embed emotions into structured spaces, using classification objectives or external supervision. For instance, (Dathathri et al., 2019) and (Buechel et al., 2020) train models to map between emotion spaces. Similarly, (Wang and Zong, 2021) learns an emotion space from labeled data, shows clustering by valence, and demonstrates transferability across datasets. However, in all of these works, the emotion space is imposed or supervised, not emergent.
A growing line of work probes how pretrained models encode emotion. (Hollinsworth et al., 2024) show that valence is linearly embedded in contextual states, while (Zhang et al., 2023) find arousal and dominance less separable, though their analysis depends on encoder-only models and fixed affective lexicons. In contrast, we study decoder-only LLMs, seeking to recover emergent emotional structure directly from hidden-state geometry rather than imposing a psychological model.
Other studies have shown that LLMs exhibit strong zero-shot emotion classification performance across languages (Bianchi et al., 2022), though subsequent work notes that language-specific tuning is sometimes necessary for culturally grounded affect (De Bruyne et al., 2022). These findings suggest that emotion representations are at least partially transferable across linguistic domains—a hypothesis we test more directly through geometric alignment and projection-based analysis in Section 4.
@@ 3 Methods
To understand how emotions are represented in LLMs, a variety of tools were used. This section outlines those methods and their theoretical grounding. Empirical findings from these analyses are presented in Sections 4 and 4.3.
@@
Centered-SVD. We build on prior work showing that LLM hidden states lie on low-dimensional manifolds where semantic and syntactic properties are linearly recoverable (Aghajanyan et al., 2020; Hu et al., 2022; Lizzo and Heck, 2025). To isolate emotion-relevant subspaces, we apply singular value decomposition (SVD) to hidden state activations.
Emotion-relevant subspaces are extracted using centered singular value decomposition (SVD) on sentence-level hidden states. For each input, token activations are mean-pooled to obtain a single sentence representation; these vectors are stacked, centered, and decomposed via SVD to obtain orthogonal directions of variation. When emotional content is the dominant structured difference across the inputs, the leading components should capture the primary affective axes in the model’s representation space and form the basis for the alignment, probing, and causal manipulation analyses described in later sections.
To interpret the semantic content of the principal components derived from the centered SVD, the relative ordering of emotion centroids are examined along each component. Component polarity is standardized by flipping signs when necessary to ensure consistent orientation across layers. This procedure yields a stable ranking of emotions along each axis, enabling cross-layer comparison and semantic interpretation of the leading components.
For qualitative visualization, we embed the mean-pooled hidden-state centroids into two dimensions using t-SNE, applied to the same activation representations used for SVD. This embedding provides a coarse, nonlinear view of cluster structure and complements the linear analyses conducted in the emotional subspace.
@@
Space Alignment. We first assess cross-dataset similarity through centroid cosine similarity, measuring how consistently each emotion’s direction aligns with its counterpart in the synthetic manifold; for each emotion, we compute the cosine between its dataset-specific centroid vector and the corresponding centroid in the synthetic space. Prior work has shown that latent spaces arising from related tasks often exhibit similar internal geometry, with relationships between them approximately rigid or linear up to rescaling and rotation (Moschella et al., 2023). While some approaches lift these spaces into anchor-relative representations to handle isometric variance, recent work demonstrates that direct alignment via linear or rigid transformations is often sufficient and easier to apply in practice (Lähner and Moeller, 2024). Following this approach, we use linear regression to align the emotional subspace derived from synthetic data with that derived from human-authored emotion datasets, allowing us to test whether the synthetic manifold reflects transferable emotional encodings or merely generation artifacts. $W^{*}$ is obtained by a least-squares fit over paired real–synthetic sentence-level activations.
$W^{*}=\arg\min_{W}\,\|\,YW-X\,\|_{F}^{2},$
The resulting MSE between $X$ and $YW^{*}$ summarizes how well the synthetic manifold can be recovered from the real space. To characterize the learned transformation, we report its Frobenius norm (overall magnitude) and spectral flatness (isotropic versus anisotropic scaling), which help distinguish genuine geometric compatibility from alignment achieved through highly directional distortion.
To evaluate how consistently emotional structure is preserved across datasets, we complement the cosine-similarity and regression-based alignment metrics with a set of high-dimensional geometry measures that quantify relational distortion between emotional spaces. Whereas cosine similarity evaluates directional agreement and regression error captures how well one space can be linearly mapped onto another, these distortion and stress metrics assess the internal geometry itself, that is, how faithfully pairwise relationships among emotions are preserved after alignment. They reveal whether two spaces share not only similar directions but also comparable distance structure, providing a stricter and more global test of representational equivalence.
Geometry Preservation Metrics. To evaluate how emotional structure is preserved across datasets, geometric deformation is quantified between emotional spaces using both classical stress measures and distortion metrics. All geometry-preservation metrics are computed over pairs of activation vectors from different datasets corresponding to the same emotion, comparing their pairwise distances in each dataset’s latent space to the distances in the synthetic latent space. Stress scores quantify aggregate embedding discrepancy, while distortion metrics assess whether those discrepancies arise from uniform rescaling or more uneven, directionally biased deformation.
Stress-1, Stress-2, and Sammon stress measure how well one distance matrix can be embedded into another. Stress-1 computes the mean-squared discrepancy between distances, Stress-2 captures the same discrepancy without square-root normalization, and Sammon stress reweights errors by the inverse of the original distance, emphasizing preservation of local structure. Pairwise distances are defined by $D^{(X)}_{ij}=\|x_{i}-x_{j}\|_{2}$ and $D^{(Y)}_{ij}=\|y_{i}-y_{j}\|_{2}$. Stress-2 is then computed as
$\frac{\sum_{i<j}(D^{(H)}_{ij}-D^{(L)}_{ij})^{2}}{\sum_{i<j}(D^{(H)}_{ij})^{2}}$ (1)
Classical thresholds for low stress (e.g., Stress-1 < 0.1) were developed for low-dimensional maps (Kruskal, 1964) and do not directly apply to high-dimensional hidden-state spaces, so we interpret these values across models. Stress-1 and Sammon stress appear in the appendix.
[…]
$\rho_{ij}=\frac{D^{(Y)}_{ij}}{D^{(X)}_{ij}+\varepsilon}$ (2)
$\ell_{2}$-distortion captures the squared deviation of stretch ratios from a single global scale, reflecting uneven expansion or contraction (ideal=0). $\sigma$-distortion measures the variance of normalized stretch ratios and therefore reflects the consistency of distortion across pairs (ideal=0). Together, the stress and distortion metrics distinguish cases where emotional spaces differ only by global rescaling from cases exhibiting more heterogeneous or anisotropic deformation. Only average distortion is reported in the main text; the remaining metrics appear in the appendix.
Layer-Level Distortion Analysis. Because distortion can vary substantially across depth, especially in transformer models, we compute distortion metrics at every sublayer and quantify the percentage of layers that exhibit high distortion for each dataset–model pair. Datasets exceeding a fixed distortion threshold (set from the distribution of values across all models and datasets) are marked as “high-distortion.” Reporting the fraction of affected layers provides a more sensitive measure of representational fragility than averaging distortion over depth, which can obscure localized failure points.
Effect of Dimensionality Reduction. When comparing emotional spaces after projecting into the 50D synthetic subspace, dimensionality reduction inevitably increases measured distortion. We therefore evaluate both the full-space and 50D cases to assess whether models with already coherent emotional structure maintain that structure under compression, and whether models with fragile geometry degrade further.
Linear Probing. We train linear classifiers on activations projected into the synthetic emotional subspace and evaluate them on human-written datasets. This tests whether emotional information remains linearly recoverable after projection, complementing the geometric alignment metrics with a measure of functional decodability.
ML-AURA. ML-AURA quantifies how selectively a neuron responds to a specific concept by framing each neuron as a threshold-based detector (Suau et al., 2024). For a labeled dataset $D$, each neuron’s output is summarized per example using the maximum activation across tokens. These scalar responses are then ranked and evaluated using the area under the precision-recall curve, comparing neuron output against the presence or absence of the target concept. Neurons with high AUC-PR are designated as “experts” for that concept.
In our adaptation, the concepts are emotion categories. We apply ML-AURA in a one-vs-all setup for each emotion, scoring each neuron by how well it distinguishes a target emotion from all others. For each layer, we compute the proportion of neurons whose one-vs-all AUROC exceeds 0.9, treating these as emotion-selective units.
4 Experiments
@@ 4.1 Datasets
To extract emotional representations, the synthetic dataset of (Reichman et al., 2025) is used, in which neutral sentences are rewritten into multiple emotions so that emotional variation becomes the principal structured difference across samples. This synthetic corpus is used solely to obtain clean, maximally polarized affect directions; all downstream evaluations, alignments, and probes are performed exclusively on human-written datasets.
The universality of emotional representations in LLMs are evaluated using eight diverse datasets, each offering explicit categorical emotion labels; datasets restricted to polarity or star ratings were excluded as too coarse. The collection spans languages, modalities, and styles: Go-Emotions contains English Reddit comments (Demszky et al., 2020), CARER covers English tweets (Saravia et al., 2018), SemEval-2007 Task 14 focuses on English news headlines (Strapparava and Mihalcea, 2007), EmoEvent includes English and Spanish tweets (Plaza-del-Arco et al., 2020), Emotions in Drama consists of German plays from the eighteenth and nineteenth centuries (Dennerlein et al., 2023), Bhaav offers Hindi short stories (Kumar et al., 2019), MultiEmotions-It provides Italian YouTube and Facebook comments (Sprugnoli and others, 2020), and EmoTextToKids features French journalistic and encyclopedic texts written for children (Étienne et al., 2024). Appendix B analyzes the structural and stylistic contents of each of these datasets.